Use the Henderson–Hasselbalch equation to calculate the pH of each solution

a solution that is 12.5 g of HF and 24.5 g of NaF in 125 mL of solution Express your answer using two decimal places.
Answer #1

Volume of the solution = 125mL = 0.125L

mass of HF= 12.5 grams

molar mass of HF = 20.01gram/mole

number of moles = mass/volume = 12.5/20.01 = 0.625 moles

concentration of HF = number of moles/volume = 0.625/0.125 =5.0M

[HF] = 5.0M

mass of NaF = 24.5 grasm

molar mass of NaF = 41.99 gram/mole

number of moles of NaF = 24.5/41.99 = 0.583 moles

concentration of NaF = 0.583/0.125 = 4.664 M

[NaF] = 4.664 M

Ka of HF = 6.6x10^-4

-log(Ka) = -log(6.6x10^-4)

PKa = 3.18

PH=PKa +log[NaF]/[HF]

PH = 3.18 +log(4.664/5.0]

PH = 3.149

PH = 3.15